OpenStudy (anonymous):

hi everyone. I am interested in evaluating the integral of an integral, using integration by parts. the problem is : dg/dx = integral(F(x,g)dt) (from a to x). it is written somewhere that g(x) = integral((x-t)F(x,g)dt) (from a to x) + c Thanks in advance.

OpenStudy (anonymous):

so you're just integrating twice?

OpenStudy (anonymous):

or are you doing a double integral?

OpenStudy (anonymous):

the problem is : dg/dx = integral(F(x,g)dt) (from a to x). it is written somewhere that g(x) = integral((x-t)F(x,g)dt) (from a to x) + c \[\frac{ dg }{ dx }=\int\limits_{a}^{x}F(x,g(t))dt\] solution is \[g(x)= c + \int\limits_{a}^{x}(x-t)F(x,g(t))dt\]

OpenStudy (anonymous):

I appreciate any answer to steps of the above solution.